OpenStudy (desmarie):

Rewrite using a negative exponent. 5/x^2

OpenStudy (desmarie):

@johnweldon1993 @mathmate

OpenStudy (mathmate):

hint: one of the laws of exponents: \(\Large a^{-n}=\frac{1}{a^n}\)

OpenStudy (desmarie):

I know that but....... would I do x^-2/5-1

OpenStudy (mathmate):

`I know that but....... would I do x^-2/5-1` Here are some examples to help you: http://www.purplemath.com/modules/exponent2.htm

OpenStudy (mathmate):

5 is a constant, and is in its simplest form.

OpenStudy (desmarie):

so what am I suppose to do with the 5?? it has to be a simple answer...not in fraction form

OpenStudy (desmarie):

@mathmate

OpenStudy (mathmate):

Wonder if you read the example to which I sent you a link. I'll give you another example: \(32/x^7=32\times (1/x^7)=32 \times x^{-7}=32x^{-7}\)

OpenStudy (desmarie):

sorry...I meant to tell you that I cant see the images.....but I did read it....before I asked for help I tried looking it up online and I clicked that website but again I could not see the images.......thanks tho I got it